If "\theta" is an angle in the third quadrant, in which quadrant does the point (sin"\theta" cos"\theta" , sin"\theta" +cos"\theta" ) lie?


1
Expert's answer
2021-10-01T00:01:20-0400

Given that "\theta" is in the third quadrant.

"\therefore sin\theta" is negative and "cos\theta" is also negative.

So, "sin\theta.cos\theta" is positive and "sin\theta+cos\theta" is negative.

"\therefore x-coordinate" is positive and "y-coordinate" is negative.

So, (sin"\theta" cos"\theta" , sin"\theta" +cos"\theta" ) lies in fourth quadrant.